# Can a calibration metric be both testable and actionable?

### v1

@@ C.2 Hardness of asymptotic ECE calibration
$\mathcal{A}:\cup_{n\geq 0}(\mathcal{X}\times[0,1])^{n}\rightarrow\{\text{ measurable functions }f:\mathcal{X}\rightarrow[0,1]\}.$ (71)
@@
$\limsup_{n\rightarrow\infty}\sup_{P}\mathbb{E}_{P^{n}}[\Delta_{\operatorname{ECE}}(f_{n})]=0,$ (72)
@@
We proceed by turning the problem into a form for which we can apply empirical process theory.
$|\widehat{\Delta}_{\operatorname{Cutoff}}(f)-\Delta_{\operatorname{Cutoff}}(f)|$ (73)
$=|\sup_{I:\text{ interval}}|\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|-\sup_{I^{\prime}:\text{ interval}}|\mathbb{E}_{n}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I^{\prime}\right\}}}}]||$ (74)
$\leq\sup_{I:\text{ interval}}||\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|-|\mathbb{E}_{n}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]||$ (75)
$\leq\sup_{I:\text{ interval}}|\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]-\mathbb{E}_{n}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|$ (76)
The first inequality follows from Lemma C.3, and the second inequality holds because of the triangle inequality.
@@
$\mathbb{P}\{|\widehat{\Delta}_{\operatorname{Cutoff}}(f)-\Delta_{\operatorname{Cutoff}}(f)|\leq 20/\sqrt{n}+\epsilon\}\geq 1-\exp(-n\epsilon^{2}/2)$ (78)
@@
$|\widehat{\Delta}_{\operatorname{Cutoff}}(f)-\Delta_{\operatorname{Cutoff}}(f)|\leq\frac{20+\sqrt{2\log(1/\delta)}}{\sqrt{n}}$ (79)
@@
$\mathbb{E}\left[\sup_{\begin{subarray}{c}f\in[-1,1]^{n}\\ \sum_{i=1}^{n=1}|f_{i+1}-f_{i}|\leq M\end{subarray}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f_{i}\right]\leq\frac{2M+1}{\sqrt{n}}$ (80)
@@
    $\sum_{i=1}^{n}\epsilon_{i}f_{i}$ $=f_{1}\sum_{i=1}^{n}\epsilon_{i}+\sum_{i=2}^{n}\epsilon_{i}(f_{i}-f_{1})$ (81)
    $=f_{1}\sum_{i=1}^{n}\epsilon_{i}+\sum_{j=2}^{n}\epsilon_{i}\sum_{j=1}^{i-1}(f_{j+1}-f_{j})$ (82)
    $=f_{1}\sum_{i=1}^{n}\epsilon_{i}+\sum_{j=1}^{n-1}(f_{j+1}-f_{j})\sum_{i=j+1}^{n}\epsilon_{i}$ (83)
@@
    $\sup_{\begin{subarray}{c}f\in[-1,1]^{n}\\ \sum_{i=1}^{n=1}|f_{i+1}-f_{i}|\leq M\end{subarray}}\sum_{i=1}^{n}\epsilon_{i}f_{i}\leq\left|\sum_{i=1}^{n}\epsilon_{i}\right|+M\max_{1\leq j\leq n-1}\left|\sum_{i=j+1}^{n}\epsilon_{i}\right|$ (84)
@@
    $\mathbb{E}\left|\sum_{i=1}^{n}\epsilon_{i}\right|\leq\ \sqrt{\mathbb{E}\left[\left(\sum_{i=1}^{n}\epsilon_{i}\right)^{2}\right]}=\sqrt{n}$ (85)
@@
    $\mathbb{E}\left[\max_{1\leq j\leq n-1}\left|\sum_{i=j+1}^{n}\epsilon_{i}\right|\right]$ $=\mathbb{E}\left[\max_{1\leq j\leq n-1}\left|\sum_{i=1}^{j}\epsilon_{i}\right|\right]$ (86)
[…]
    $=2\mathbb{E}\left|\sum_{i=1}^{n-1}\epsilon_{i}\right|$ (89)
    $=2\sqrt{n-1}$ (90)
@@
    $\mathbb{E}\left[\sup_{\begin{subarray}{c}f\in[-1,1]^{n}\\ \sum_{i=1}^{n=1}|f_{i+1}-f_{i}|\leq M\end{subarray}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f_{i}\right]\leq\frac{\sqrt{n}+2M\sqrt{n-1}}{n}\leq\frac{2M+1}{\sqrt{n}}$ (91)
Lemma C.3.
$|\sup_{g\in\mathcal{G}}\mathbb{E}g(Z)-\sup_{\tilde{g}\in\mathcal{G}}\mathbb{E}_{n}\tilde{g}(Z)|\leq\sup_{g\in\mathcal{G}}|\mathbb{E}g(Z)-\mathbb{E}_{n}g(Z)|$ (92)
  • Proof.
    Suppose not. Suppose $|\sup_{g\in\mathcal{G}}\mathbb{E}g(Z)-\sup_{\tilde{g}\in\mathcal{G}}\mathbb{E}_{n}\tilde{g}(Z)|>\sup_{g\in\mathcal{G}}|\mathbb{E}g(Z)-\mathbb{E}_{n}g(Z)|$
    Therefore, $\exists c\in\mathbb{R}$ such that
    $\sup_{g\in\mathcal{G}}|\mathbb{E}g(Z)-\mathbb{E}_{n}g(Z)|\leq c$ (93)
    while
    $|\sup_{g\in\mathcal{G}}\mathbb{E}g(Z)-\sup_{\tilde{g}\in\mathcal{G}}\mathbb{E}_{n}\tilde{g}(Z)|>c.$ (94)
    This implies
    $\mathbb{E}[g(Z)]\leq\mathbb{E}_{n}[g(Z)]+c$ (95)
    and
    $\mathbb{E}_{n}[g(Z)]\leq\mathbb{E}[g(Z)]+c.$ (96)
    Therefore,
    $\sup_{g\in\mathcal{G}}\mathbb{E}g(Z)\leq\mathbb{\sup}_{g\in\mathcal{G}}\mathbb{E}_{n}g(Z)+c$ (97)
    and
    $\sup_{g\in\mathcal{G}}\mathbb{E}_{n}g(Z)\leq\mathbb{\sup}_{g\in\mathcal{G}}\mathbb{E}g(Z)+c.$ (98)
    This implies
    $|\sup_{g\in\mathcal{G}}\mathbb{E}g(Z)-\sup_{\tilde{g}\in\mathcal{G}}\mathbb{E}_{n}\tilde{g}(Z)|\leq c$ (99)
    which is a contradiction. ∎
@@ Appendix D Post-hoc calibration: proofs and counter-examples
$\begin{split}\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq&\sup_{I:\text{ interval}}|\mathbb{E}[Y{{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}\mid h_{n}]-\mathbb{E}_{n}[h_{n}(f(X)){{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}]|\\ &+\sup_{I:\text{ interval}}|\mathbb{E}_{n}[h_{n}(f(X)){{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}]-\mathbb{E}[h_{n}(f(X)){{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}\mid h_{n}]|\end{split}$ (100)
@@
$\begin{split}\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq&\sup_{I:\text{ interval}}|\mathbb{E}[Y{{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}\mid h_{n}]-\mathbb{E}_{n}[Y{{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}]|\\ &+\sup_{I:\text{ interval}}|\mathbb{E}_{n}[h_{n}(f(X)){{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}]-\mathbb{E}[h_{n}(f(X)){{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}\mid h_{n}]|\end{split}$ (101)
@@
$\begin{split}\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq&\sup_{I:\text{ interval}}|\mathbb{E}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}\mid h_{n}]-\mathbb{E}_{n}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|\\ &+\sup_{I:\text{ interval}}|\mathbb{E}_{n}[h_{n}(f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]-\mathbb{E}[h_{n}(f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}\mid h_{n}]|\end{split}$ (102)
@@
$\begin{split}\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq&\sup_{I:\text{ interval}}|\mathbb{E}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}\mid h_{n}]-\mathbb{E}_{n}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|\\ &+\sup_{I:\text{ interval}}|\mathbb{E}_{n}[{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]-\mathbb{E}[{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}\mid h_{n}]|\end{split}$ (103)
@@
$\begin{split}\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq&\sup_{I:\text{ interval}}|\mathbb{E}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]-\mathbb{E}_{n}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|\\ &+\sup_{I:\text{ interval}}|\mathbb{E}_{n}[{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]-\mathbb{E}[{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|.\end{split}$ (104)
Using a similar Rademacher complexity argument from Proposition 4.1, we know with probability at least $21-\delta/$ 2 that each term individually is bounded above by $\frac{20+\sqrt{2\log(2/\delta)}}{\sqrt{n}}$ and $\frac{10+\sqrt{2\log(2/\delta)}}{\sqrt{n}}$, respectively. Note that the high-probability bound for the first term is higher than the second due to the contraction mapping argument doubling the Rademacher complexity.
@@
$\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq\frac{30+2\sqrt{2\log(2/\delta)}}{\sqrt{n}}$ (105)
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{f}(X_{i}){{\mathbf{1}_{\left\{\hat{f}(X_{i})\in A\right\}}}}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}{{\mathbf{1}_{\left\{\hat{f}(X_{i})\in A\right\}}}}$ (106)
@@
$|\mathbb{E}_{P}[g(X)]-\mathbb{E}_{Q}[g(X)]|\leq\sup_{\text{Interval }I\subseteq\mathbb{R}}|P(I)-Q(I)|$ (107)
@@
    $|\mathbb{E}_{P}[g(X)]-\mathbb{E}_{Q}[g(X)]|$ $=\left|\int_{\mathbb{R}}g(x)(dP-dQ)(x)\right|$ (108)
[…]
    $\leq\sup_{t\in[0,1]}\left|\int_{\mathbb{R}}{{\mathbf{1}_{\left\{g(x)\geq t\right\}}}}(dP-dQ)(x)\right|$ (111)
@@
    $\sup_{t\in[0,1]}\left|\int_{\mathbb{R}}{{\mathbf{1}_{\left\{g(x)\geq t\right\}}}}(dP-dQ)(x)\right|\leq\sup_{\text{Interval }I\subseteq\mathbb{R}}|P(I)-Q(I)|$ (113)

### v2

@@ C.2 Hardness of asymptotic ECE calibration
$\mathcal{A}:\cup_{n\geq 0}(\mathcal{X}\times[0,1])^{n}\rightarrow\{\text{ measurable functions }f:\mathcal{X}\rightarrow[0,1]\}.$
@@
$\limsup_{n\rightarrow\infty}\sup_{P}\mathbb{E}_{P^{n}}[\Delta_{\operatorname{ECE}}(f_{n})]=0,$
@@
  • Proof of Proposition 4.1.
    We proceed by turning the problem into a form for which we can apply empirical process theory. For notational brevity, we adopt the notation that $\mathbb{E}_{n}[g(X,Y)]:=\frac{1}{n}\sum_{i=1}^{n}g(X_{i},Y_{i})$, where $g$ is an arbitrary function.
    Observe that
    $|\widehat{\Delta}_{\operatorname{Cutoff}}(f)-\Delta_{\operatorname{Cutoff}}(f)|$
    $=\left|\sup_{I:\text{ interval}}\left|\mathbb{E}\left[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}\right]\right|-\sup_{I^{\prime}:\text{ interval}}\left|\mathbb{E}_{n}\left[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I^{\prime}\right\}}}}\right]\right|\right|$
    $\leq\sup_{I:\text{ interval}}||\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|-|\mathbb{E}_{n}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]||$
    $\leq\sup_{I:\text{ interval}}|\mathbb{E}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]-\mathbb{E}_{n}[(Y-f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|,$
    where the first inequality follows from the triangle inequality of suprema.
@@
    $\mathbb{P}\{|\widehat{\Delta}_{\operatorname{Cutoff}}(f)-\Delta_{\operatorname{Cutoff}}(f)|\leq 20/\sqrt{n}+\epsilon\}\geq 1-\exp(-n\epsilon^{2}/2).$
@@
    $|\widehat{\Delta}_{\operatorname{Cutoff}}(f)-\Delta_{\operatorname{Cutoff}}(f)|\leq\frac{20+\sqrt{2\log(1/\delta)}}{\sqrt{n}}$
@@
$\mathbb{E}\left[\sup_{\begin{subarray}{c}f\in[-1,1]^{n}\\ \sum_{i=1}^{n=1}|f_{i+1}-f_{i}|\leq M\end{subarray}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f_{i}\right]\leq\frac{2M+1}{\sqrt{n}}.$
@@
    $\sum_{i=1}^{n}\epsilon_{i}f_{i}$ $=f_{1}\sum_{i=1}^{n}\epsilon_{i}+\sum_{i=2}^{n}\epsilon_{i}(f_{i}-f_{1})$
    $=f_{1}\sum_{i=1}^{n}\epsilon_{i}+\sum_{i=2}^{n}\epsilon_{i}\sum_{j=1}^{i-1}(f_{j+1}-f_{j})$
    $=f_{1}\sum_{i=1}^{n}\epsilon_{i}+\sum_{j=1}^{n-1}(f_{j+1}-f_{j})\sum_{i=j+1}^{n}\epsilon_{i}.$
@@
    $\sup_{\begin{subarray}{c}f\in[-1,1]^{n}\\ \sum_{i=1}^{n=1}|f_{i+1}-f_{i}|\leq M\end{subarray}}\sum_{i=1}^{n}\epsilon_{i}f_{i}\leq\left|\sum_{i=1}^{n}\epsilon_{i}\right|+M\max_{1\leq j\leq n-1}\left|\sum_{i=j+1}^{n}\epsilon_{i}\right|.$
@@
    $\mathbb{E}\left|\sum_{i=1}^{n}\epsilon_{i}\right|\leq\ \sqrt{\mathbb{E}\left[\left(\sum_{i=1}^{n}\epsilon_{i}\right)^{2}\right]}=\sqrt{n}.$
@@
    $\mathbb{E}\left[\max_{1\leq j\leq n-1}\left|\sum_{i=j+1}^{n}\epsilon_{i}\right|\right]$ $=\mathbb{E}\left[\max_{1\leq j\leq n-1}\left|\sum_{i=1}^{j}\epsilon_{i}\right|\right]$
    $=\int_{0}^{\infty}\mathbb{P}\left\{\max_{1\leq j\leq n-1}\left|\sum_{i=1}^{j}\epsilon_{i}\right|\geq t\right\}dt$
    $\leq 2\int_{0}^{\infty}\mathbb{P}\left\{\left|\sum_{i=1}^{n-1}\epsilon_{i}\right|\geq t\right\}dt$
    $=2\mathbb{E}\left|\sum_{i=1}^{n-1}\epsilon_{i}\right|$
    $=2\sqrt{n-1}.$
@@
    $\mathbb{E}\left[\sup_{\begin{subarray}{c}f\in[-1,1]^{n}\\ \sum_{i=1}^{n=1}|f_{i+1}-f_{i}|\leq M\end{subarray}}\frac{1}{n}\sum_{i=1}^{n}\epsilon_{i}f_{i}\right]\leq\frac{\sqrt{n}+2M\sqrt{n-1}}{n}\leq\frac{2M+1}{\sqrt{n}}.$
@@ Appendix D Post-hoc calibration: proofs and counter-examples
    $\begin{split}\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq&\sup_{I:\text{ interval}}|\mathbb{E}[Y{{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}\mid h_{n}]-\mathbb{E}_{n}[h_{n}(f(X)){{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}]|\\ &+\sup_{I:\text{ interval}}|\mathbb{E}_{n}[h_{n}(f(X)){{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}]-\mathbb{E}[h_{n}(f(X)){{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}\mid h_{n}]|.\end{split}$
@@
    $\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq$ $\sup_{I:\text{ interval}}|\mathbb{E}[Y{{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}\mid h_{n}]-\mathbb{E}_{n}[Y{{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}]|$ (5)
    $+\sup_{I:\text{ interval}}|\mathbb{E}_{n}[h_{n}(f(X)){{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}]-\mathbb{E}[h_{n}(f(X)){{\mathbf{1}_{\left\{h_{n}(f(X))\in I\right\}}}}\mid h_{n}]|.$
@@
$\begin{split}\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq&\sup_{I:\text{ interval}}|\mathbb{E}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}\mid h_{n}]-\mathbb{E}_{n}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|\\ &+\sup_{I:\text{ interval}}|\mathbb{E}_{n}[h_{n}(f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]-\mathbb{E}[h_{n}(f(X)){{\mathbf{1}_{\left\{f(X)\in I\right\}}}}\mid h_{n}]|.\end{split}$
@@
$\begin{split}\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq&\sup_{I:\text{ interval}}|\mathbb{E}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}\mid h_{n}]-\mathbb{E}_{n}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|\\ &+\sup_{I:\text{ interval}}|\mathbb{E}_{n}[{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]-\mathbb{E}[{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}\mid h_{n}]|.\end{split}$
@@
$\begin{split}\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq&\sup_{I:\text{ interval}}|\mathbb{E}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]-\mathbb{E}_{n}[Y{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|\\ &+\sup_{I:\text{ interval}}|\mathbb{E}_{n}[{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]-\mathbb{E}[{{\mathbf{1}_{\left\{f(X)\in I\right\}}}}]|.\end{split}$
Using a similar Rademacher complexity argument from Proposition 4.1, we know with probability at least $1-\delta/2$ that each term individually is bounded above by $\frac{20+\sqrt{2\log(2/\delta)}}{\sqrt{n}}$ and $\frac{10+\sqrt{2\log(2/\delta)}}{\sqrt{n}}$, respectively. Note that the high-probability bound for the first term is higher than the second due to the contraction mapping argument doubling the Rademacher complexity.
@@
$\Delta_{\operatorname{Cutoff}}^{P}(h_{n}\circ f)\leq\frac{30+2\sqrt{2\log(2/\delta)}}{\sqrt{n}},$
@@
$\frac{1}{n}\sum_{i=1}^{n}\hat{f}(X_{i}){{\mathbf{1}_{\left\{\hat{f}(X_{i})\in A\right\}}}}=\frac{1}{n}\sum_{i=1}^{n}Y_{i}{{\mathbf{1}_{\left\{\hat{f}(X_{i})\in A\right\}}}}.$
@@
$|\mathbb{E}_{P}[g(X)]-\mathbb{E}_{Q}[g(X)]|\leq\sup_{\text{Interval }I\subseteq\mathbb{R}}|P(I)-Q(I)|.$
@@
    Observe that
    $|\mathbb{E}_{P}[g(X)]-\mathbb{E}_{Q}[g(X)]|$ $=\left|\int_{\mathbb{R}}g(x)(dP-dQ)(x)\right|$
    $=\left|\int_{\mathbb{R}}\int_{0}^{1}{{\mathbf{1}_{\left\{t\leq g(x)\right\}}}}dt(dP-dQ)(x)\right|$
    $=\left|\int_{0}^{1}\int_{\mathbb{R}}{{\mathbf{1}_{\left\{g(x)\geq t\right\}}}}(dP-dQ)(x)dt\right|$
    $\leq\sup_{t\in[0,1]}\left|\int_{\mathbb{R}}{{\mathbf{1}_{\left\{g(x)\geq t\right\}}}}(dP-dQ)(x)\right|.$
@@
    $\sup_{t\in[0,1]}\left|\int_{\mathbb{R}}{{\mathbf{1}_{\left\{g(x)\geq t\right\}}}}(dP-dQ)(x)\right|\leq\sup_{\text{Interval }I\subseteq\mathbb{R}}|P(I)-Q(I)|.$